# SADDE: Semi-supervised Anomaly Detection with Dependable Explanations

### v1

@@ 5. Related Work
Interpreters. Numerous studies have been conducted to interpret machine learning models and achieved promising results, which can be roughly categorized into two groups, i.e., local/model-agnostic (Ribeiro et al., 2016a, 2018; Guidotti et al., 2018a; Lundberg, 2017; Guo et al., 2018; Fong et al., 2019) and global interpreters (Fisher et al., 2019; Kauffmann et al., 2024). The former interprets individual predictions for a given machine learning model while the latter attempts to explain model predictions of a cluster. Local/model-agnostic interpreters are mostly perturbation-based, Gradient-based, and others. Perturbation-based approaches (Fong et al., 2019; Fong and Vedaldi, 2017; Yang et al., 2021) perturb data features and check the variation of model predictions to identify the most important ones. Gradient-based approaches (Shrikumar et al., 2017b; Selvaraju et al., 2017) back-propagate gradients through the model to evaluate each feature’s sensitivity. Others treat the target machine learning model as a black-box (Ribeiro et al., 2016b, 2018). Techniques like LIME (Ribeiro et al., 2016b), LIMNA (Guo et al., 2018), and SHAP (Lundberg and Lee, 2017) attempt to utilize a linear model to approximate the decision boundary of the input, and utilize it to find the most influential features. Global interpreters like (Fisher et al., 2019) explain model prediction by randomly permuting specific feature values of a cluster and observer model predictions. The authors of (Kauffmann et al., 2024) developed an interpreter for explaining clustering algorithms by formulating clustering decisions as being functionally equivalent neural networks.
@@
Few studies aim to interpret unsupervised models such as COIN (Liu et al., 2017b) and CADE (Yang et al., 2021) which are used as baselines in our experiments. In addition to them, most of other works directly use the interpreters proposed for supervised learning to unsupervised learning applications.
@@
The work of (Fan et al., 2020) demonstrated that the explanation results provided by existing interpreters lack consensus, resulting in confusion regarding their trustworthiness. There are a few studies that examine the reliability of existing interpreters (Yeh et al., 2019; Camburu et al., 2019; Fan et al., 2020), for example, (Yeh et al., 2019) introduced two metrics to evaluate interpreters, i.e., sensitivity and infidelity. An off-the-shelf evaluation test is proposed for post-hoc interpreters in view of feature selection in (Camburu et al., 2019). (Fan et al., 2020) introduced three evaluation metrics to measure interpreters’ stability, effectiveness, and robustness. The work of (Warnecke et al., 2020) developed criteria for evaluating and comparing different interpreters, including both general properties like accuracy and security-related ones, such as robustness, efficiency, and completeness. Unfortunately, to our best knowledge, there is no existing work exploring how to improve the reliability of interpretation results.
@@
Semi-supervised Learning. A recent surge of interest has focused on anomaly detection with semi-supervised learning, which attempts to assign high-confidence pseudo-labels to unlabeled data. There is a wide range of research on semi-supervised learning, which generally falls into two categories, i.e., single-stage-based and multi-stage-based approaches. Some single-stage-based semi-supervised learning techniques improve the confidence of pseudo labels based on thresholds (Li et al., 2024; Min et al., 2024), and they usually set high thresholds to unlabeled samples to prevent the occurrence of incorrect pseudo labels, or contrastive learning (Fini et al., 2023; Wei and Gan, 2023) that combines combine pseudo labeling and consistency regularization to encourages similar predictions between two different views of an instance to improve the robustness of the model. Unlike single-stage-based semi-supervised learning techniques, multi-stage-based semi-supervised learning approaches, such as (Sun et al., 2020a, b), explore the most confident pseudo labels by using multiple models and improve their confidence by aligning the predictions from each stage.
6. Conclusion
@@
In this paper, we propose SADDE, a general framework aiming to interpret and improve semi-supervised anomaly detection performance. It incorporates two key techniques KAM and ToS: (1) KAM provides reliable, human-understandable, stable, robust, and fast interpretations for semi-supervised systems; (2) ToS makes it possible to automatically learn from large unlabeled datasets. By applying and evaluating Adaptive NAD over two classical network anomaly detection datasets, we demonstrate that SADDE can provide fast and accurate anomaly detection results as well as reliable, human-understandable, stable, robust, and fast interpretations.

### v4

@@ 6 Related Work
Interpreters. Numerous studies have been conducted to interpret machine learning models and achieved promising results, which can be roughly categorized into two groups, i.e., local/model-agnostic [ribeiro2016should, ribeiro2018anchors, lundberg2017unified_shap, guo2018_lemna, fong2019understanding] and global interpreters [fisher2019all, 9817459]. The former interprets individual predictions for a given machine learning model, while the latter attempts to explain model predictions of a cluster. Local/model-agnostic interpreters are mostly perturbation-based, Gradient-based, and others. Perturbation-based approaches [fong2019understanding, fong2017interpretable, yang2021cade] perturb data features and check the variation of model predictions to identify the most important ones. Gradient-based approaches [shrikumar2017learning_deeplift] back-propagate gradients through the model to evaluate each feature’s sensitivity. Others treat the target machine learning model as a black-box [ribeiro2016should_lime, ribeiro2018anchors]. Techniques like LIME [ribeiro2016should_lime], LIMNA [guo2018_lemna], and SHAP [lundberg2017unified_shap] attempt to utilize a linear model to approximate the decision boundary of the input, and utilize it to find the most influential features. Global interpreters like [fisher2019all] explain model predictions by randomly permuting specific feature values of a cluster and observer model predictions. The authors of [9817459] developed an interpreter for explaining clustering algorithms by formulating clustering decisions as being functionally equivalent neural networks.
@@
Few studies aim to interpret unsupervised models such as COIN [liu2017contextual_coin] and CADE [yang2021cade], which are used as baselines in our experiments. In addition to them, most of the other works directly use the interpreters proposed for supervised learning to unsupervised learning applications.
@@
There are a few studies that examine the reliability of existing interpreters [yeh2019fidelity, fan2020can], for example, the authors of [yeh2019fidelity] introduced two metrics to evaluate interpreters, i.e., sensitivity and infidelity. Fan et al. [fan2020can] introduced three evaluation metrics to measure interpreters’ stability, effectiveness, and robustness. The work of [9230374] developed criteria for evaluating and comparing different interpreters, including both general properties like accuracy and security-related ones, such as robustness, efficiency, and completeness. Unfortunately, to our best knowledge, there is no existing work exploring how to improve the reliability of interpretation results.
@@
Semi-supervised Learning. A recent surge of interest has focused on anomaly detection with semi-supervised learning, which attempts to assign high-confidence pseudo-labels to unlabeled data. There is a wide range of research on semi-supervised learning, which generally falls into two categories, i.e., single-stage-based and multi-stage-based approaches. Some single-stage-based semi-supervised learning techniques improve the confidence of pseudo labels based on thresholds [li2024instantT, min2024leveraging_LVM], and they usually set high thresholds to unlabeled samples to prevent the occurrence of incorrect pseudo labels, or contrastive learning [Fini_2023_CVPR_clustering, wei2023towards_ACR] that combines combine pseudo labeling and consistency regularization to encourages similar predictions between two different views of an instance to improve the robustness of the model. Unlike single-stage-based semi-supervised learning techniques, multi-stage-based semi-supervised learning approaches, such as [sun2020multi_M3S, sun2020multi], explore the most confident pseudo labels by using multiple models and improve their confidence by aligning the predictions from each stage.
7 Conclusion
@@
In this paper, we propose AnomalyAID, a general framework aiming to interpret and improve semi-supervised anomaly detection performance. It incorporates two key techniques, KAM and ToS: (1) KAM provides reliable interpretations for semi-supervised systems; (2) ToS makes it possible to automatically learn from large unlabeled datasets. By applying and evaluating Adaptive NAD over two classical network anomaly detection datasets, we demonstrate that AnomalyAID can provide fast and accurate anomaly detection results as well as reliable interpretations. As part of future work, researchers can further investigate the performance of AnomalyAID on various anomaly detection scenarios with different model combinations under extreme data distributions.
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (Grant No. 62406215) and in part by the supported by the ”Fundamental Research Funds for the Central Universities”.